OpenStudy (compassionate):

Let f(x) = 2x - 6. Solve f-1(x) when x = 2. So, how would I do this? I'm finding an inverse.

OpenStudy (rational):



OpenStudy (akashdeepdeb):

Yes, you have to find the inverse here. What is the inverse? It is just another graph g(x) [inverse of f(x)] which is symmetrical to the original graph f(x) about the x = y line. Here, f(x) = 2x-6 Let y = 2x-6 Now, represent, 'x' in terms of y. y + 6 = 2x x = \(\frac{y+6}{2}\) Now, switch the variables. y = \(\frac{x+6}{2}\) \[f^{-1} (x) = \frac{x+6}{2}\]

OpenStudy (rational):